OpenStudy (anonymous):

Which inequality expresses all the solutions to 4 ≤ 8 − 2x < 6?

OpenStudy (anonymous):

−3 < x ≤ 2 1 < x ≤ 2 x ≤ − 3 or x ≥ 2 x ≤ 1 or x ≥ 2

OpenStudy (astrophysics):

Start off by subtracting 8 from both sides what do you get

OpenStudy (anonymous):

B

OpenStudy (astrophysics):

...

OpenStudy (anonymous):

4 ≤ − 2x < -2?

OpenStudy (astrophysics):

Close it should be -4 right?

OpenStudy (astrophysics):

\[-4 \le -2x < -2\] right

OpenStudy (anonymous):

oh i subtracted from the wrong side lol ya

OpenStudy (astrophysics):

Haha it's cool, ok now we can just divide -2 on both sides but note that when we divide by a negative all the signs flip

OpenStudy (anonymous):

\[-2 \ge x < -1\]

OpenStudy (anonymous):

like this?

OpenStudy (astrophysics):

Close, when you divide a negative by a negative it's positive right?

OpenStudy (anonymous):

is its 1 not -1?

OpenStudy (astrophysics):

It should be \[2 \ge x > 1\]

OpenStudy (astrophysics):

Which should be written as \[1 < x \le 2\]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thats what i said

OpenStudy (astrophysics):

But you gave a direct answer without explaining the process which has no meaning here

OpenStudy (anonymous):

ty astro your the best

OpenStudy (astrophysics):

Anytime